what is the answer to 5:3 (ratio) if three equals 615?

Answer by DrBeeee(684)

Given:
(1) 5/3 = x/615
Solve for x by multiplying both sides of (1) by 615 to get
(2) 615*5/3 = 615*x/615 or
(3) 205*5 = 1*x or
(4) x = 1025
Check this with (1).
Is (5/3 = 1025/615)?
Is (5*615 = 3*1025)?
Is (3075 = 3075)? Yes
Answer: If 3 is equivalent to 615 in the ratio 5:3, then the 5 is equivalent to 1025.